OpenStudy (anonymous):

Find the number of sides of a regular polygon if one interior angle is 60 degrees. A.) 3 B.) 4 C.) 5 D.) 6

OpenStudy (anonymous):

I'm not so sure how to solve this, can someone please take me through this! :)

OpenStudy (mathstudent55):

Do you know the formula for the sum of the measures of the interior angles of a polygon?

OpenStudy (anonymous):

isn't it S= 180(n-2) ?

OpenStudy (mathstudent55):

It is (n - 2)180, where n = number of sides. Since you have a regular polygon, the measure of each interior angle is the sum of the measures of the angles divided by the number of anlges.

jimthompson5910 (jim_thompson5910):

If k is the interior angle, then we can use this formula k = 180(n-2)/n plug in k = 60 to get 60 = 180(n-2)/n 60n = 180(n-2) 60n = 180n - 360 I'll let you finish in solving for n

OpenStudy (mathstudent55):

That gives you ((n - 2)180) / n, but this equals 60 in your problem, so you get ((n - 2)180)/n = 60 Now solve the equation for n

OpenStudy (anonymous):

I got three when I divided 120 from 360, does that seem right? :O

jimthompson5910 (jim_thompson5910):

n = 3 is correct

jimthompson5910 (jim_thompson5910):

here's what it looks like 

OpenStudy (anonymous):

that makes a lot more sense!! thank you so much!! :)

jimthompson5910 (jim_thompson5910):

yw